# Accurate Identification of Communication Between Multiple Interacting Neural Populations

### v1

@@ 1 Introduction
In this work, we introduce Multi-Region Latent Factor Analysis via Dynamical Systems (MR-LFADS), a multi-region communication model that directly addresses all of the challenges outlined above. MR-LFADS is a probabilistic model that represents each recorded region with a distinct set of stacked recurrent neural networks (RNNs) that capture the region’s potentially nonlinear and nonstationary population dynamics. MR-LFADS represents communication between observed regions and inputs from unobserved regions as disentangled sets of latent variables. Structured information bottlenecks encourage the model to infer inputs from unobserved regions only when their effects cannot be explained by communication among the recorded regions. MR-LFADS infers single-trial initial conditions and time-varying inputs that together account for trial-to-trial variability in the recorded activity. This automatic inference of inputs eliminates the need to manually specify input signals, thereby avoiding strong, difficult-to-validate assumptions about how external signals influence each region. Finally, MR-LFADS constrains communication to originate from model-reconstructed neural activity, rather than from more-flexible latent representations—a design choice that, as we will show, enables more accurate inference of communication without sacrificing the quality of data reconstruction.
@@ 2 Related Work
Existing communication models can be broadly categorized as either static or dynamic. Static methods predict each timestep of neural activity in a target region from a corresponding timestep of activity in one or more source regions and then interpret predictive source activity as inter-regional communication (Kaufman et al., 2014; Perich et al., 2018; Ruff & Cohen, 2019; Veuthey et al., 2020). Reduced-rank regression (RRR) is a prominent static technique, which models target-region activity as a low-rank linear function of the source-region activity (Semedo et al., 2019; Steinmetz et al., 2019; MacDowell et al., 2025). While static methods are straightforward to fit and interpret, they are typically limited to capturing instantaneous, linear dependencies. Consequently, they do not readily account for nonlinear or nonstationary relationships, or temporal structure that may arise due to neural population dynamics (Vyas et al., 2020).
@@
While these existing approaches to communication modeling address some of the challenges outlined in Section 1, none, to our knowledge, address all four challenges. In particular, none of these existing methods support inferring inputs from unobserved brain regions. This functionality is crucial because inputs from unobserved regions might influence the target region’s population dynamics and modes of communication. Some approaches attempt to account for unobserved inputs by explicitly providing task-related signals as inputs to each region or by removing condition averages and modeling the residual single-trial neural activity. However, these manual strategies impose strong assumptions about the content and targets of input signals. As we will show, misspecifying such inputs risks confounding inferred population dynamics and communication.
@@
Figure 1: MR-LFADS architecture. (a) Adaptation of single-region LFADS with Poisson outputs. (b) MR-LFADS with $N$ regions. KL penalties from SR-LFADS in panel (a) are included in MR-LFADS, but are omitted in the diagram for clarity.
@@ 3 Multi-Region LFADS (MR-LFADS)
Notation. All time-indexed variables and parameters are also indexed by trial, though we omit trial indices for notational simplicity. We use $t_{1}$:$t_{2}$ to denote the inclusive sequence of integers $\{t_{1},t_{1}+1,\dots,t_{2}\}$. We use $W^{i}(x):=W^{i}x+b^{i}$ to denote an affine transformation with weights $W^{i}$ and offsets $b^{i}$.
@@
$P(x_{t}^{i}\mid g_{0}^{i},u_{1:t}^{i},\{m_{1:t}^{j\rightarrow i}\}_{j\neq i})=\mathcal{N}(r_{t}^{i},\Sigma_{r,t}^{i})\vspace{-3pt}$ (4)
@@
$\mu^{j\rightarrow i}_{m,t}$ $={W}^{j\rightarrow i}_{\mu_{m}}(r_{t}^{j})$ (8)
$\Sigma^{j\rightarrow i}_{m,t}$ $=\mbox{diag}\Big{(}\exp\big{(}W_{\sigma_{m}}^{j\rightarrow i}(r_{t}^{j})\big{)}\Big{)}$
@@ 4 Results I: Synthetic Multi-Region Datasets
To benchmark MR-LFADS, we compare it to three established communication modeling techniques—RRR, mp-srSLDS, and MR-SDS (see Section 2). To demonstrate the importance of specific MR-LFADS design features, we also compare against ablated MR-LFADS variants that selectively exclude those features. We focus evaluation on each method’s ability to recover a causal model of each DGN—including both the pathways and content of inter-regional communication—given a multi-region dataset generated by the DGN.
@@
To assess communication content, we quantify how well the model-inferred messages capture the information in the ground truth messages. Specifically, we apply linear regression to predict the ground truth messages $m_{t}^{j\rightarrow i}$ from the inferred messages $\mu_{m,t}^{j\rightarrow i}$ (Eq. 7), and we report the coefficient of determination, denoted $R^{2}(\mu_{m}^{j\rightarrow i},m^{j\rightarrow i})$. For MR-LFADS, we use an analogous procedure to compare model-inferred inputs $\mu_{u,t}^{i}$ (Eq. 10) to ground truth external inputs. This comparison is not applicable to RRR, mp-srSLDS, or MR-SDS, as these methods do not infer inputs from unobserved regions. Additional details on experiments, model hyperparameters, and evaluation metrics are provided in Appendix B-D.
4.1 Experiment 1: Inputs to Modeled Regions
@@
Figure 2: Experiment 1. (a) Left: DGN setup. Each region, area $A^{i}$, receives a private stimulus $s_{t}^{i}$ (red) and communicates a two-step delayed version $s_{t-2}^{i}$ (blue) to a downstream region. Each region is trained to recall (green) the last five time steps of its private stimulus and its received communication. Right: Potential incorrect communication model capable of accurately reconstructing these synthetic neural data. (b) $R^{2}$ scores for data reconstruction. (c) Left: Cosine similarity between inferred effectomes and ground truth. Right: Example fitted models. Arrow weights indicate average message norms across trials and time. (d) $R^{2}$ of linear prediction of ground truth stimulus inputs (with time lag $\in\{0,2\}$) from inferred inputs. (e) $R^{2}$ of linear prediction of ground truth messages (with time lag $\in\{2,4\}$) from inferred messages.
@@ 4.2 Experiment 2: Data-Constrained Communication
Figure 3: Experiment 2. (a) Left: DGN setup, with stimulus (red), communication (blue), trained readouts (green), and computations (I: identity; $\int$: integration). Right: Potential failure mode for a learned model. (b) $R^{2}$ scores for data reconstruction. (c) Left: Cosine similarity between inferred effectomes and ground truth. Right: Example fitted models. (d) $R^{2}$ of linear prediction of ground truth input $s$ to region $P$, from inferred inputs. (e) Left: $R^{2}$ when predicting ground truth messages $m^{P\rightarrow D}$$=$$s$ from inferred messages, $\mu_{m}^{P\rightarrow D}$. Right: $R^{2}$ when predicting the decision variable $d$ from $\mu_{m}^{P\rightarrow D}$, indicating mislocalization of integration.
@@
4.3 Experiment 3: Random Multi-Region Networks
Figure 4: Experiment 3. (a) Top: Each DGN is configured with a random number of areas ($3$ or $4$), random inter-regional connectivity, and is trained on a randomly selected task. Bottom: Example task setup. On each trial, the DGNs received a noisy fixation stimulus $s_{\text{fix},t}$ and two noisy task stimuli, represented in polar coordinates as $s_{t}^{k}=(a_{t}^{(k)},\theta_{t}^{(k)})$ for $k\in\{1,2\}$. DGNs were trained to process these inputs into task-dependent outputs: a response angle $\theta^{\text{resp}}$ and a task-dependent saccadic eye movement. Here, multisensory decision-making is depicted as an example task. When the fixation cue disappears, the output area $A^{N}$ must saccade and report the $\theta^{(i)}$ value corresponding to the stimulus with larger $a^{(i)}$. (b) $R^{2}$ scores for data reconstruction. (c) Cosine similarity ($S_{\text{cos}}$) of inferred effectomes relative to ground truth, compared to that of MR-LFADS(R), with $\Delta S_{\text{cos}}=S_{\text{cos}}^{\text{model}}-S_{\text{cos}}^{\text{\mbox{MR-LFADS(R)}}}$. One-tailed t-test p-values: $p=0.00016$, $0.12$, $0.006$, $0.0$, $0.01$. NS: not significant. (d) Left: $R^{2}$ scores, relative to MR-LFADS(R), for linear prediction of ground truth messages from inferred messages. One-tailed t-test p-values: $p=0.0$, $0.0$, $0.001$, $0.0$, $0.14$. Right: $R^{2}$ scores, relative to MR-LFADS(R), for linear prediction of inferred messages from ground truth messages. One-tailed t-test p-values: $p=0.0004$, $0.0$, $0.08$, $0.0$, $0.0$. (e) Inferred effectomes from models fit to datasets from a three-region DGN (top) and a four-region DGN (bottom). We chose the datasets on which MR-LFADS(R) achieved its median $S_{\text{cos}}$ scores (across the 35 generated datasets). Arrow thicknesses indicate relative message norms.
@@ 5 Results II: Multi-Region Electrophysiology
Figure 5: MR-LFADS(R) applied to multi-region, high-density electrophysiology. (a) Mice receive a high- or low-tone auditory stimulus (“sample”) and respond by licking left or right (“response”). (b) MR-LFADS single-trial predicted firing rates in held-out control trials (blue), recorded spike times (black vertical ticks), and smoothed, binned spike counts (black; causal, exponential filter) for example neurons in each modeled brain region. (c) Condition-averaged smoothed spike counts of example neurons in control and photoinhibition trials (left-hit condition). Shaded regions indicate the standard deviation across trials. (d) Photoinhibition-related changes in population recordings, as observed experimentally (left) and as predicted by MR-LFADS (right). (e) Cosine similarity of inferred effectomes across models with different random initializations (seeds). (f) Message norms inferred by MR-LFADS(R) for all connections across example seeds $k$. (g) Correlation of inferred message norms across pairs of seeds $(k,l)$.
@@ 6 Discussion
Understanding how brain regions interact to support distributed computation requires communication models that can disentangle inter-regional communication from local population dynamics and region-specific inputs from unrecorded regions. In this work, we identified critical failure modes that limit existing communication models—including misidentification due to manually specified external inputs and mislocalization of neural dynamics due to overly flexible communication architectures. We introduced MR-LFADS, a communication model specifically designed to mitigate such failures through three key design features: (1) automatic inference of region-specific inputs from unobserved sources, (2) data-constrained communication inferred from reconstructed firing rates, and (3) structured regularization that promotes disentangling and prevents inferring inputs from unobserved sources when the same information can be obtained via communication from observed regions. These features discourage the model from learning spurious solutions that explain the data but misrepresent inter-regional interactions.
@@
Another set of MR-LFADS HPs (the $\beta$’s from Section 3) control the model’s preference to infer communication rather than inputs from unobserved regions whenever possible. While the settings we chose were effective in this study, different choices may be needed in other scenarios—especially when an unrecorded region directly drives multiple recorded regions. This classical correlation-versus-causation confound likely cannot be resolved when fitting to passively observed (rather than causally perturbed) neural activity. Although we found that MR-LFADS can recover more causal structure than existing methods, its greatest utility may lie in generating hypotheses about communication and guiding targeted causal circuit manipulations to test them.
@@ A.1 Multi-Region LFADS (MR-LFADS)
Reconstruction validation. During training, we held out $10\%$ of the neurons in each real or synthetic brain region, and we used the activity of these held-out neurons to evaluate model fit. A separate linear decoder is trained post hoc to predict held-out neural activity from the inferred factors. Importantly, this decoder is not trained end-to-end with MR-LFADS, ensuring that held-out neurons do not influence model learning. Fit quality is quantified using the $R^{2}$ score between predicted and actual held-out activity.
@@
Table S1: Key hyperparameters for MR-LFADS models. Experiment 4 refers to applications to multi-region electrophysiology.
@@
T 190 Total time used for inferring inferred inputs
$\tau$ 10 Total time used for inferring the initial condition
@@
$\beta_{u}$ start epoch 50 Epoch at which $\beta_{u}$ starts increasing from 00
@@
$\beta_{m}$ start epoch 150 Epoch at which $\beta_{m}$ starts increasing from 00
@@
$\alpha$ start epoch 0 Epoch at which $\alpha$ starts increasing from 00
@@
$N^{i}_{\text{fac}}$ $N^{i}_{\text{Neu}}$ Factor size

### v2

@@ 1 Introduction
In this work, we introduce Multi-Region Latent Factor Analysis via Dynamical Systems (MR-LFADS), a multi-region communication model that directly addresses all of the challenges outlined above. MR-LFADS is a probabilistic model that represents each recorded region with a distinct set of stacked recurrent neural networks (RNNs) that capture the region’s potentially nonlinear and nonstationary population dynamics. MR-LFADS represents communication between observed regions and inputs from unobserved regions as disentangled sets of latent variables. Structured information bottlenecks encourage the model to infer inputs from unobserved regions only when their effects cannot be explained by communication among the recorded regions. MR-LFADS infers single-trial initial conditions and time-varying inputs that together account for trial-to-trial variability in the recorded activity. This automatic inference of inputs eliminates the need to manually specify input signals, thereby avoiding strong, difficult-to-validate assumptions about how external signals influence each region. Finally, MR-LFADS constrains communication to originate from model-reconstructed neural activity, rather than from more flexible latent representations—a design choice that, as we will show, enables more accurate inference of communication without sacrificing the quality of data reconstruction.
@@ 2 Related Work
Existing communication models can be broadly categorized as either static or dynamic. Static methods predict each timestep of neural activity in a target region from a corresponding timestep of activity in one or more source regions and then interpret predictive source activity as inter-regional communication (Kaufman et al., 2014; Perich et al., 2018; Ruff & Cohen, 2019; Veuthey et al., 2020). Reduced-rank regression (RRR) is a prominent static technique that models target-region activity as a low-rank linear function of the source-region activity (Semedo et al., 2019; MacDowell et al., 2025). While static methods are straightforward to fit and interpret, they are typically limited to capturing instantaneous, linear dependencies. Consequently, they do not readily account for nonlinear or nonstationary relationships, or temporal structure that may arise due to neural population dynamics (Vyas et al., 2020).
@@
While these existing approaches to communication modeling address some of the challenges outlined in Section 1, none, to our knowledge, address all four challenges. In particular, none of these existing methods support inferring inputs from unobserved brain regions. This functionality is crucial because inputs from unobserved regions might influence the target region’s population dynamics and modes of communication. Some approaches attempt to account for unobserved inputs by explicitly providing task-related signals as inputs to each region or by removing condition averages and modeling the residual single-trial neural activity. However, these manual strategies impose strong assumptions about the content and targets of input signals. As we will show, misspecifying such inputs risks confounding inferred population dynamics and communication, leading to models that accurately reconstruct neural activity through incorrect mechanisms.
@@
Figure 1: MR-LFADS architecture. (a) Single-region LFADS, as adapted for this work. (b) MR-LFADS with NN regions. KL penalties from SR-LFADS in panel (a) are included in MR-LFADS, but are omitted in the diagram for clarity.
@@ 3 Multi-Region LFADS (MR-LFADS)
Notation. All time-indexed variables and parameters are also indexed by trial, though we omit trial indices for notational simplicity. We use $t_{1}$$t_{2}$ to denote the inclusive sequence of integers $\{t_{1},t_{1}+1,\dots,t_{2}\}$. We use $W^{i}(x)\hskip-2.5pt:=\hskip-1.99997ptW^{i}x+b^{i}$ to denote an affine transformation with weights $W^{i}$ and offsets $b^{i}$.
@@
$P(x_{t}^{i}\mid g_{0}^{i},u_{1:t}^{i},\{m_{1:t}^{j\rightarrow i}\}_{j\neq i})=\mathcal{N}(r_{t}^{i},\Sigma_{r,t}^{i})\vskip-3.0pt$ (4)
@@
$\hskip-10.0pt\begin{aligned} \mu^{j\rightarrow i}_{m,t}&={W}^{j\rightarrow i}_{\mu_{m}}(r_{t}^{j})\\ \Sigma^{j\rightarrow i}_{m,t}&=\mbox{diag}\Big{(}\exp\big{(}W_{\sigma_{m}}^{j\rightarrow i}(r_{t}^{j})\big{)}\Big{)}\end{aligned}$ (8)
@@ 4 Results I: Synthetic Multi-Region Datasets
To benchmark MR-LFADS, we compare it to three established communication modeling techniques—RRR (see Appendix A.2), mp-srSLDS, and MR-SDS (see Section 2). To demonstrate the importance of specific MR-LFADS design features, we also compare against ablated MR-LFADS variants that selectively exclude those features. We focus evaluation on each method’s ability to recover a causal model of each DGN—including both the pathways and content of inter-regional communication—given a multi-region dataset generated by the DGN.
@@
To assess communication content, we quantify how well the model-inferred messages capture the information in the ground truth messages. Specifically, we apply linear regression to predict the ground truth messages $m_{t}^{j\rightarrow i}$ from the inferred messages $\mu_{m,t}^{j\rightarrow i}$ (Eq. 7), and we report a cross-validated coefficient of determination, denoted $R^{2}(\mu_{m}^{j\rightarrow i},m^{j\rightarrow i})$. For MR-LFADS, we use an analogous procedure to compare model-inferred inputs $\mu_{u,t}^{i}$ (Eq. 10) to ground truth external inputs. This comparison is not applicable to RRR, mp-srSLDS, or MR-SDS, as these methods do not infer inputs from unobserved regions. Additional details on experiments, model hyperparameters, and evaluation metrics are provided in Appendix B-D.
4.1 Experiment 1: Inferring Unobserved Inputs
@@
Figure 2: Experiment 1. (a) Left: DGN setup. Each region, area $A^{i}$, receives a private stimulus $s_{t}^{i}$ (red) and communicates a two-step delayed version $s_{t-2}^{i}$ (blue) to a downstream region. Each region is trained to recall (green) the last five time steps of its private stimulus and its received communication. Right: Potential incorrect communication model capable of accurately reconstructing these synthetic neural data. (b) $R^{2}$ scores for data reconstruction. Box plots describe distributions of values across $10$ models fit from distinct random initializations (seeds). Boxes represent the interquartile range (IQR), and whiskers extend to the most extreme points within $1.5$ IQR from the quartiles. (c) Left: Cosine similarity between inferred effectomes and ground truth. Right: Example fitted models. Color intensity indicates the relative message norm, computed by concatenating all multidimensional messages across trials and time, taking the 2-norm, then normalizing across communication channels. $s^{1:3}$ indicates the ground truth input. (d) $R^{2}$ of linear prediction of ground truth stimulus inputs (with time lag $\in\{0,2\}$) from inferred inputs. (e) $R^{2}$ of linear prediction of ground truth messages (with time lag $\in\{2,4\}$) from inferred messages.
@@ 4.2 Experiment 2: Data-Constrained Communication
Figure 3: Experiment 2. (a) Left: DGN setup, with stimulus (red), communication (blue), trained readouts (green), and computations (I: identity; $\int$: integration). Right: Potential failure mode for a learned model. (b) $R^{2}$ scores for data reconstruction. (c) Left: Cosine similarity between inferred effectomes and ground truth. Right: Example fitted models. Color intensity indicates relative message norm. (d) $R^{2}$ of linear prediction of ground truth input $s$ to region $P$, from inferred inputs. (e) Left: $R^{2}$ when predicting ground truth messages $m^{P\rightarrow D}$==$s$ from inferred messages, $\mu_{m}^{P\rightarrow D}$. Right: $R^{2}$ when predicting the decision variable $d$ from $\mu_{m}^{P\rightarrow D}$, indicating mislocalization of integration.
@@
Figure 4: Experiment 3. (a) Top: Each DGN is configured with a random number of areas (33 or 44), random inter-regional connectivity, and is trained on a randomly selected task. Bottom: Example task setup. On each trial, the DGNs received a noisy fixation stimulus $s_{\text{fix},t}$ and two noisy task stimuli, represented in polar coordinates as $s_{t}^{k}=(a_{t}^{(k)},\theta_{t}^{(k)})$ for $k\in\{1,2\}$. DGNs were trained to process these inputs into task-dependent outputs: a response angle $\theta^{\text{resp}}$ and a task-dependent saccadic eye movement. Here, multisensory decision-making is depicted as an example task. When the fixation cue disappears, the output area $A^{N}$ must saccade and report the $\theta^{(i)}$ value corresponding to the stimulus with larger $a^{(i)}$. (b) $R^{2}$ scores for data reconstruction. Box plots describe distributions of values across $35$ DGN fits. (c) Cosine similarity ($S_{\text{cos}}$) of inferred effectomes relative to ground truth, compared to that of MR-LFADS(R), with $\Delta S_{\text{cos}}=S_{\text{cos}}^{\text{model}}-S_{\text{cos}}^{\text{\mbox{MR-LFADS(R)}}}$. One-tailed t-test p-values: $p=0.00016$, $0.12$, $0.006$, $0.0$, $0.01$. NS: not significant. (d) Left: $R^{2}$ scores, relative to MR-LFADS(R), for linear prediction of ground truth messages from inferred messages. One-tailed t-test p-values: $p=0.0$, $0.0$, $0.001$, $0.0$, $0.14$. Right: $R^{2}$ scores, relative to MR-LFADS(R), for linear prediction of inferred messages from ground truth messages. One-tailed t-test p-values: $p=0.0004$, $0.0$, $0.08$, $0.0$, $0.0$. (e) Inferred effectomes from models fit to datasets from a three-region DGN (top) and a four-region DGN (bottom). We chose the datasets on which MR-LFADS(R) achieved its median $S_{\text{cos}}$ scores (across the 35 generated datasets). Color intensity indicates relative message norm.
@@
4.3 Experiment 3: Generalization Across Random Multi-Region Networks
@@ 5 Results II: Multi-Region Electrophysiology
Figure 5: MR-LFADS(R) applied to multi-region, high-density electrophysiology. (a) Mice receive a high- or low-tone auditory stimulus (“sample”) and respond by licking left or right (“response”). (b) MR-LFADS single-trial predicted firing rates in held-out control trials (blue), recorded spike times (black vertical ticks), and smoothed, binned spike counts (black; causal, exponential filter) for example neurons in each modeled brain region. (c) Condition-averaged smoothed spike counts of example neurons in control and photoinhibition trials (left-hit condition). Shaded regions indicate the standard deviation across trials. (d) Photoinhibition-related changes in population recordings, as observed experimentally (left) and as predicted by MR-LFADS (right). (e) Cosine similarity of inferred effectomes across models with different random initializations (seeds). (f) Message norms inferred by MR-LFADS(R) for all connections across example seeds $k$. (g) Correlation of inferred message norms across pairs of seeds $(k,l)$. Box plots describe distributions of values across 55 models fit from distinct random initializations.
@@ 6 Discussion
Understanding how brain regions interact to support distributed computation requires communication models that can disentangle inter-regional communication from local population dynamics, accounting for region-specific inputs from unrecorded regions. In this work, we identified critical failure modes that limit existing communication models—including misidentification due to manually specified external inputs and mislocalization of neural dynamics due to overly flexible communication architectures. We introduced MR-LFADS, a communication model specifically designed to mitigate such failures through three key design features: (1) automatic inference of region-specific inputs from unobserved sources, (2) data-constrained communication inferred from reconstructed firing rates, and (3) structured regularization that promotes disentangling and prevents inferring inputs from unobserved sources when the same information can be obtained via communication from observed regions. These features discourage the model from learning spurious solutions that explain the data but misrepresent inter-regional interactions. While MR-LFADS is one concrete implementation, the three design principles are more general, and alternative architectures—for example, different formulations of the local-region dynamical systems—may also succeed if they incorporate these same principles.
@@
Another set of MR-LFADS HPs (the $\beta$’s from Section 3) control the model’s preference to infer communication rather than inputs from unobserved regions whenever possible. While the settings we chose were effective in this study, different choices may be needed in other scenarios—especially when an unrecorded region directly drives multiple recorded regions. This classical correlation–causation confound cannot be resolved from passively observed neural activity alone. While MR-LFADS recovers more causal structure than existing methods, its greatest value may be in generating hypotheses about communication and motivating targeted causal circuit perturbations to test them.
@@ A.1 Multi-Region LFADS (MR-LFADS)
Reconstruction validation. During training, we held out $10\%$ of the neurons in each real or synthetic brain region for validation. After training MR-LFADS, we fit a separate linear decoder post hoc on the training trials, regressing the inferred factors onto the held-out neurons’ activity. Since the decoder is not trained end-to-end with MR-LFADS, held-out neurons do not influence model fitting. Goodness of fit is reported as the $R^{2}$ between predicted and observed held-out-neuron activity on validation trials.
@@
Table S1: Key hyperparameters for MR-LFADS models. Experiment 4 refers to applications to multi-region electrophysiology data.
@@
T 190 Total time steps used for inferring inferred inputs
$\tau$ 10 Total time steps used for inferring the initial condition
@@
$\beta_{u}$ start epoch 50 Epoch at which $\beta_{u}$ starts increasing from 0
@@
$\beta_{m}$ start epoch 150 Epoch at which $\beta_{m}$ starts increasing from 0
@@
$\alpha$ start epoch 0 Epoch at which $\alpha$ starts increasing from 0
@@
$^{i}_{\text{fac}} NN$$^{i}_{\text{Neu}}$ Factor size